Europium metal crystallizes in a body–centered lattice with the unit cell measuring 4.58 

A° along each edge. Determine the volume of a unit cell.


1
Expert's answer
2021-06-07T03:15:44-0400

A body-centred lattice forms a cubic unit cell. As a result, the volume of a cube equals:

V = a3

where a is the edge length of the cube. 

From here, the volume of a unit cell equals:

V = a3 = (4.58 A°)3 = (4.58 × 10-10 m)3 = 9.6 × 10-29 m3


Answer: 9.6 × 10-29 m3